OpenStudy (anonymous):

A subway train travels from one station to the next in 2 min. Its distance, in km, from the first station after t minutes is s(t) = t^2 - 1/3 t^3 . At what time will the train have a velocity of 0.5km/min?

OpenStudy (anonymous):

\[s(t) = t ^{2}- 1/3 t ^{3} \] so i know velocity = displacement / time then what should i do?

OpenStudy (anonymous):

\[v(t)=\frac{d}{dt}s(t)=-t^2+2t=0.5\] http://www.wolframalpha.com/input/?i=2x^2-4x%2B1+%3D+0